# The distribution of interior branch lengths in phylogenetic trees under constant-rate birth-death models

### v1

@@ Abstract
Phylogenetic trees represent the evolutionary relationships between extant lineages, omitting extinct or non-sampled lineages. Extending the work of Stadler and collaborators, this paper focuses on the distribution of branch lengths in phylogenetic trees arising under a constant-rate birth-death model. We derive branch length distributions of interior branches with and without random sampling of individuals of the extant population. We establish that branches connected to the tree leaves and interior branches behave very differently under sampling: pendant branches get longer as the sampling probability reaches zero, whereas the distribution of interior branches quickly reaches an asymptotic state.
@@ I Introduction
The reconstructed evolutionary process is central to understanding evolutionary relationships and dynamics. This reconstruction is based on extant genetic data and yields a phylogenetic tree, which represents the evolutionary relationships among a set of species, genes, or cells. Crucially, all lineages that eventually went extinct (or were not sampled) are missing from this tree. This affects the statistics of the reconstructed tree: Under a stochastic birth-death process, species go extinct at a certain rate (or cells die, in a cellular system). Looking at the reconstructed tree, only speciation events are seen. To compare empirical phylogenetic trees with models of evolution, the link between an evolutionary process and the reconstructed process it induces must be understood.
@@
The reconstructed evolutionary process has been studied extensively, with early work of Nee, May, and Harvey Nee et al. (1994) giving the probability density for each speciation event in a reconstructed tree from a birth-death process with constant rates at a particular time. Such a birth-death model describes a population of species, individuals, or cells, which divide independently at a fixed rate $\lambda$ and die independently at a fixed rate $\mu$. (For species, birth events describe the emergence of a new species, whereas death events correspond to extinctions.) A constant-rate birth-death model serves as a simple model of neutral evolution.
@@
The work by Nee, May, and Harvey Nee et al. (1994) has been extended by Stadler and collaborators to trees reconstructed from individuals that were sampled from a population Stadler (2009) and to populations conditioned on a particular size Gernhard (2008). The statistics of speciation events have been used to infer trees and their characteristics Stadler (2013b); MacPherson et al. (2022), and is the cornerstone of widely-used computational packages like Beast Bouckaert et al. (2019) and TreeTime Sagulenko et al. (2018). Many extensions of the basic constant-rate birth-death process have been developed, for instance, to account for rates that are variable through time or differ across the tree (selection) Stadler (2013b).
@@
In Stadler and Steel (2012); Mooers et al. (2012) Stadler, Steel and collaborators focus on a particular aspect of tree statistics, namely the distribution of branch lengths (in calendar time). They calculate the probability density function for the length of the so-called pendant branches (the lowest branches leading to the leaves of the tree) for trees reconstructed from a constant-rate birth-death process. For the particular case of a Yule process (a birth-death process with the rate of death set to zero), they also give the probability density function for the length of interior nodes.
@@
In this paper, we revisit the problem of the distribution of branch lengths with a particular focus on the lengths of interior branches and the distribution of branch lengths under sampling of extant species. In Section II, we recapitulate the relevant properties of a constant-rate birth-death process and in Section III we introduce a simple but effective approach to calculate pendant branch lengths, recovering the results of Stadler and Steel Stadler and Steel (2012). We generalise these results to the lengths of interior branches: in Section IV branches one and two levels above a pendant branch are discussed, and in Section V we turn to generic branches picked randomly in the interior of a phylogenetic tree. In Section VI, we briefly discuss the statistics of the number of generations along branches, and in Section VII we generalise these results to a scenario where only a finite fraction of the extant population is sampled. We find that pendant and interior branch lengths behave very differently under sampling: while pendant branches become longer and grow without limit as the sampling probability approaches zero, the distribution of interior branch lengths quickly reaches an asymptotic limit as the sampling probability decreases.
@@ II Dynamics of a birth-death process
A powerful alternative to the master equation 1 is a master equation for the probability that a population that started with a single individual at time zero has population size one at time $t$ Maddison et al. (2007); Stadler (2009). Here, we generalise this approach to the probability that in a population that started with a single individual at time zero, at time $t$ there is one and only one individual that has surviving offspring at a later time $T$. This probability turns out to be a key tool to calculate the distribution of branch lengths.
Specifically, $\tilde{p}^{(T)}_{1}(t)$ denotes the probability that at time $t$ a clade which started at time zero contains one individual which has at least one extant descendant (including potentially itself) at time $T$, and all other individuals alive at time $t$ have no extant descendants at time $T$ (again including themselves). The master equation for $\tilde{p}^{(T)}_{1}(t)$ is
@@
where the first term accounts for the decay of probability due to the death of that individual, and the second term describes the probability that any additional lineage arising in a birth event fails to become extinct by the time of observation. The factor of two comes from the two offspring in such a birth event which can play this role. This is a master equation in a single probability only; the population aspect of this dynamics resides compactly in the probability $p_{0}(T-t)$ that a particular clade dies out in a time interval of length $T-t$ given by (2).
The master equation (6) is a first-order linear differential equation solved by the integrating factor $-\int_{0}^{t}\operatorname{\mathrm{d}\!}t^{\prime}\left[\mu+\lambda\left(1-2p_{0}(T-t^{\prime})\right)\right]$. Integrating to $t=T$ yields the probability of population size on at time $T$ (3). Integrating $t$ only to some intermediate time $T-\tau_{e}$ gives the probability $\tilde{p}^{(T)}_{1}(T-\tau_{e})$. (The name $\tau_{e}$ is chosen as this will later mark the end of an interior branch.)
It turns out to be convenient for the analysis of branch lengths to consider times relative to the time of observation $T$. We denote the length of the interval between the intermediate time $t$ and the observation time $T$ by $\tau_{e}=T-t$. We further define $\tau=T-\tau_{e}(=t)$ as the length of the interval between $\tau_{e}$ and the time of the clade’s origin. Specifically, we denote as $p_{1}^{(\tau_{e})}(\tau)=\tilde{p}^{(T)}_{1}(t=T-\tau_{e})$ the probability that a clade started at time $\tau_{e}+\tau$ in the past, and contains at time $\tau_{e}$ in the past only a single individual that will have extant descendants at the time of observation. Again, setting $\tau_{e}=0$ results in $p_{1}^{(\tau_{e}=0)}(\tau)\equiv p_{1}(\tau)$. For the case of $\tau_{e}>0$ the integrating factor becomes $-\int_{0}^{T-\tau_{e}}\operatorname{\mathrm{d}\!}t^{\prime}\left[\mu+\lambda\left(1-2p_{0}(T-t^{\prime})\right)\right]=-\int_{0}^{\tau}\operatorname{\mathrm{d}\!}\tau^{\prime}\left[\mu+\lambda\left(1-2p_{0}(\tau_{e}+\tau^{\prime})\right)\right]=\log\left(\frac{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}})^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{e}+\tau)})^{2}}\right)$ and we obtain
@@
This probability differs from $p_{1}(\tau)\equiv p^{(\tau_{e}=0)}_{1}(\tau)$, as for $\tau_{e}>0$ additional lineages have more time to die out than in pendant branches (specifically $\tau_{e}$ more). This will lead to a difference between the statistics of interior and pendant branches.
@@
In the following sections, we will always measure times $\tau$ relative to the time of observation, increasing as one looks further into the evolutionary past, see Figure 1
@@ III Branch length distribution: pendant branches
The results from section II can be used directly to determine the statistics of branch lengths of a phylogenetic tree. In this section, we apply them to the so-called pendant branches, which end in a leaf of the phylogenetic tree. In this way, we recover the results of Stadler and Steel Stadler and Steel (2012). In the following sections, we apply the same technique to branches in the interior of a phylogenetic tree.
A pendant branch (a branch terminating at the time of observation) of length $\tau$ was created in a birth event at time $\tau$ , where (i) one of the two offspring produces a clade consisting of a single extant member at the time of observation, and (ii) the other offspring produces a clade that did not die out by the time of observation. Figure 2A shows a simple example. Conditions (i) and (ii) are statistically independent events whose probabilities can be expressed in terms of the quantities computed above. This leaves the rate at which birth events occur per small time interval, which is given by the birth rate $\lambda$ times the population size at a particular time.
We consider the case where the number of individuals $N$ at the time of observation is known, but no further information is given, for instance, on the age of the tree. Since the population size is unknown at time $\tau\neq 0$, we sum over all possible population sizes and their probabilities $\sum_{n=1}^{\infty}np_{n}(\tau)$. Ignatieva et al. (2020) have shown that the statistics of a population of size $N$ tracked backwards in time under a birth-death process is described by a birth-death process with time running forwards and birth and death rates exchanged. The expected population size of the forward process is proportional to $e^{(\lambda-\mu)t}$. Hence the expected population size at time $\tau$ before the observation is proportional to $\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}$. The dependence on $N$ (which can be summed over) is only in the constant of proportionality.
The relative number of pendant branches of length $\tau$ per small time interval is therefore given by
@@
The factor of two at the end of the first term accounts for the emergence of two branches of length $\tau$ when both offspring only have a single surviving descendant. $p_{>1}(\tau)=\sum_{n=2}^{\infty}p_{n}(\tau)$ denotes the probability that a clade founded at time $\tau$ has more than one extant member at the time of observation. Using equations (2) and (3), we obtain the probability density function of the lengths $\tau$ of pendant branches (up to a normalising factor found to be one) as
@@
Superficially, this result looks different from the distribution of pendant branch lengths derived in Stadler and Steel Stadler and Steel (2012) and Mooers et al. Mooers et al. (2012), which in the present notation is $2\lambda p_{1}(\tau)\left(1-\frac{\lambda}{\mu}p_{0}(\tau)\right)$. The results agree, as they must, since $\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}\left(1-p_{0}(\tau)\right)=\frac{(\lambda-\mu)\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}=1-\frac{\lambda}{\mu}p_{0}(\tau)$.
@@ IV Branch length distribution: interior branches above pendant branches
Figure 2: Statistics of pendant and interior branches. (A) This schematic phylogenetic tree shows birth events (circles) and death events ($\times$). The pendant branch of length $\tau$ on the left arose from a birth event in the past (solid circle), where one offspring (shown left) has exactly one extant descendant, and the other offspring (shown right) has at least one extant descendant. This particular pendant branch has one generation between its origin at time $\tau$ measured from the time of observation in the past and the present (not counting the birth event at $\tau$). (B) A birth event (top solid node) leading to a level-two branch of length $\tau_{2}$ and a level-one (pendant) branch of length $\tau_{1}$. (C) A schematic phylogenetic tree with branches of different levels, see text. (D) The distribution of branch lengths $\tau_{2}$ of second-level branches in a population of $N=7500$ individuals (blue bars), grown in a simulation as described in the text, compared to the theoretical probability density function (11) marginalised over the pendant branch lengths $\tau_{1}$ (red line). (E) The distribution of interior branch lengths in a population of $N=7500$ individuals (blue bars), compared to the theoretical density function (14) (red line).
Since the probability of a birth event at time $\tau_{1}+\tau_{2}$ was $\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{1}+\tau_{2})}$ (up to a normalising prefactor, see below), the relative number of level-two and level-one branches of lengths $\tau_{2}$ and $\tau_{1}$ per small time intervals is thus
@@
Equation (11) gives the joint probability density function of branch lengths of pendant and level-two branches up to a normalisation factor, which again turns out to be one. Figure 2D compares this result to numerical simulations. A population with birth rate $\lambda=1$ and death rate $\mu=0.75$ was grown from a single individual to a population of $7500$ individuals (stopping when the population first reaches that number), and the phylogenetic tree was reconstructed. The lengths of level-two branches were collected in a histogram and compared to the joint probability $P_{\text{level}}(\tau_{1},\tau_{2})$ given by (11) marginalised with respect to $\tau_{1}$. We note that this simulation produces statistics that are slightly different from the statistics derived above. The simulation halts when a particular population size is reached, whereas the calculation looks at the statistics at a particular time $T$, see Hartmann et al. (2010) for the subtle distinction. At large population sizes, the fluctuations around the mean population size and, hence, the difference between these two ensembles are small.

### v2

# Branch length statistics in phylogenetic trees under constant-rate birth-death models

@@ Abstract
Phylogenetic trees represent the evolutionary relationships between extant lineages, where extinct or non-sampled lineages are omitted. Extending the work of Stadler and collaborators, this paper focuses on the branch lengths in phylogenetic trees arising under a constant-rate birth-death model. We derive branch length distributions of phylogenetic branches with and without random sampling of individuals of the extant population under two distinct statistical scenarios: a fixed age of the birth-death process and a fixed number of individuals at the time of observation. We find that branches connected to the tree leaves (pendant branches) and branches in the interior of the tree behave very differently under sampling: pendant branches get longer as the sampling probability is decreased, whereas the distribution of interior branches quickly reaches an asymptotic regime.
@@ I Introduction
The reconstructed evolutionary process is central to understanding evolutionary relationships and dynamics. This reconstruction is based on extant genetic data and yields a phylogenetic tree, which represents the evolutionary relationships among a set of species, genes, or cells. Crucially, all lineages that eventually went extinct (or were not sampled) are missing from such a tree. This affects the statistics of the reconstructed tree: Under a stochastic birth-death process, species go extinct at a specific rate (or cells die, in a cellular system). Looking at the reconstructed tree, however, only speciation events are seen. To compare empirical phylogenetic trees with models of evolution, it is crucial to understand the link between an evolutionary process and the reconstructed process.
@@
The reconstructed evolutionary process has been studied extensively, with early work of Nee, May, and Harvey Nee et al. (1997) giving the probability density in time for each speciation event in a reconstructed tree from a birth-death process with constant rates. Such a birth-death model describes a population of species, individuals, or cells, which divide independently at a fixed rate $\lambda$ and die independently at a fixed rate $\mu$. (For species, birth events describe the emergence of a new species, whereas death events correspond to extinctions.) A constant-rate birth-death model serves as a simple model of neutral evolution.
@@
Stadler and collaborators have extended the work by Nee, May, and Harvey Nee et al. (1997) to trees reconstructed from individuals that were sampled from a population Stadler (2009) and to populations conditioned on a particular size Gernhard (2008). The resulting statistics of speciation events have been used to infer trees and their characteristics Stadler (2013b); MacPherson et al. (2022), and is the cornerstone of widely-used computational packages such as BEAST Bouckaert et al. (2019) and TreeTime Sagulenko et al. (2018). Many extensions of the basic constant-rate birth-death process have been developed, for instance, to account for rates that are variable through time or vary across the tree (selection) Stadler (2013b).
@@
In Stadler and Steel (2012); Mooers et al. (2012) Stadler, Steel and collaborators focus on a particular aspect of tree statistics, namely the distribution of branch lengths (in calendar time). They calculate the probability density function for the length of the so-called pendant branches (the lowest branches leading to the leaves of the tree) for trees reconstructed from a constant-rate birth-death process. For the particular case of a Yule process (a birth-death process with the rate of death set to zero), they also give the probability density function for the length of interior branches.
In this paper, we revisit the problem of the distribution of branch lengths with a particular focus on the lengths of interior branches and the distribution of branch lengths under sampling of extant species. Two statistical scenarios need to be distinguished: In the first scenario (scenario i)), the time interval between the start of the birth-death process, starting with a single individual, and the time of observation (the end of the birth-death process) is known, but the number of extant species at the end of that process is a fluctuating random variable. In the second scenario (scenario ii)), one conditions on the number of individuals at the end of the process, but the time interval between the start and end of the process is unknown.
@@
In Section II, we recapitulate the relevant properties of a constant-rate birth-death process and in Sections III-V, we discuss a simple but effective approach to determine branch length distributions under the first scenario used by Paradis Paradis (2016). We write down concrete expressions for branch length distributions in the limit of a long-running birth-death process, focussing on pendant branches in Section III, on branches one and two levels above a pendant branch in Section IV, and on generic branches picked randomly in the interior of a phylogenetic tree in Section V. Section VI briefly discusses the statistics of the number of generations along branches. In Section VII we generalize these results to a scenario where only a finite fraction of the extant population is sampled. We find that pendant and interior branch lengths behave very differently under sampling: while pendant branches become longer and grow without limit as the sampling probability approaches zero, the distribution of interior branch lengths quickly reaches an asymptotic limit as the sampling probability decreases. In Section VIII we turn to the second scenario (conditioning on the number of extant species at the time of observation) and introduce a novel approach based on tracking the birth-death process first backwards in time from the time of observation to the emergence of a particular branch and then again forwards to the time of observation. For pendant branches, we recover the results of Stadler, Steel and collaborators Stadler and Steel (2012); Mooers et al. (2012) and generalize these results to interior branches.
@@ II Dynamics of a birth-death process
A powerful alternative to the master equation (1) is a master equation for the probability that a population that started with a single individual at time zero has population size one at time $t$ Maddison et al. (2007); Stadler (2010). A generalization of this approach is the probability that in a population that started with a single individual at time zero, at time $t$ there is one and only one individual that has surviving offspring at a later time $T$. This probability turns out to be a key tool to calculate the distribution of branch lengths.
Specifically, $\tilde{p}^{(T)}_{1}(t)$ denotes the probability that at time $t$ a clade which started at time zero contains one individual with open future fate, and all other individuals alive at time $t$ have no extant descendants at time $T$ (including themselves). The master equation for $\tilde{p}^{(T)}_{1}(t)$ is
@@
where the first term accounts for the decay of probability due to that individual’s death, and the second term describes the probability that any additional lineage arising in a birth event fails to become extinct by the time of observation. The factor of two comes from the two offspring in such a birth event which can play this role. This is a master equation in a single probability only; the population aspect of this dynamics resides compactly in the probability $p_{0}(T-t)$ that a particular clade dies out in a time interval of length $T-t$ given by (2).
The master equation (6) is a first-order linear differential equation solved by the integrating factor $-\int_{0}^{t}\operatorname{\mathrm{d}\!}t^{\prime}\left[\mu+\lambda\left(1-2p_{0}(T-t^{\prime})\right)\right]$. Integrating to $t=T$ yields the probability of population size one at time $T$ (3). Integrating $t$ only to some intermediate time $T-\tau_{e}$ gives the probability $\tilde{p}^{(T)}_{1}(T-\tau_{e})$. (The name $\tau_{e}$ is chosen as this will later mark the end of an interior branch.)
It turns out to be convenient for the analysis of branch lengths to consider times relative to the time of observation $T$. We denote the length of the interval between the intermediate time $t$ and the observation time $T$ by $\tau_{e}=T-t$. We further define $\tau=T-\tau_{e}$ as the length of the interval between $\tau_{e}$ and the time of the clade’s origin. Specifically, we denote as $p_{1}^{(\tau_{e})}(\tau)=\tilde{p}^{(T)}_{1}(t=T-\tau_{e})$ the probability that a clade started at time $\tau_{e}+\tau$ in the past, and contains at time $\tau_{e}$ in the past only a single individual that could have extant descendants at the time of observation. Again, setting $\tau_{e}=0$ results in $p_{1}^{(\tau_{e}=0)}(\tau)\equiv p_{1}(\tau)$. For the case of $\tau_{e}>0$ the integrating factor becomes $-\int_{0}^{T-\tau_{e}}\operatorname{\mathrm{d}\!}t^{\prime}\left[\mu+\lambda\left(1-2p_{0}(T-t^{\prime})\right)\right]=-\int_{0}^{\tau}\operatorname{\mathrm{d}\!}\tau^{\prime}\left[\mu+\lambda\left(1-2p_{0}(\tau_{e}+\tau^{\prime})\right)\right]=\log\left(\frac{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}})^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}}{(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{e}+\tau)})^{2}}\right)$ and we obtain
@@
This probability differs from $p_{1}(\tau)\equiv p^{(\tau_{e}=0)}_{1}(\tau)$, as for $\tau_{e}>0$ additional lineages have more time to die out than in pendant branches (specifically $\tau_{e}$ more). This will lead to a difference between the branch length statistics of interior and pendant branches. In the following sections, we will measure times $\tau$ relative to the time of observation, with $\tau$ increasing as one looks further into the evolutionary past, see Figure 1.
@@ III Branch length distribution: pendant branches
The results from section II can be used directly to determine the statistics of branch lengths of a phylogenetic tree. In this section, we apply them to the so-called pendant branches, which end in a leaf of the phylogenetic tree. We use the approach of Paradis Paradis (2016) to determine branch length distributions and write down concrete expressions for the probability densities. This section and the following are concerned with complete sampling (where the phylogenetic tree is constructed from the entire population alive at a particular time since the start of the birth-death process), in Section VII we will turn to the case where only a finite fraction of the population is available to reconstruct a tree from.
A pendant branch (a branch terminating at the time of observation) of length $\tau$ was created in a birth event at time $\tau$ , where (i) one of the two offspring produces a clade consisting of a single extant member at the time of observation, and (ii) the other offspring produces a clade that did not die out by the time of observation. Figure 2A shows a simple example. Conditions (i) and (ii) are statistically independent events, whose probabilities can be expressed in terms of the quantities computed above. This leaves the rate at which birth events occur per small time interval, which is given by the birth rate $\lambda$ times the population size at time $\tau$. Since the population size is unknown, we sum over all possible population sizes $n$ and their probabilities, yielding that the probability of a birth event at time $\tau$ as the expectation value of the population size multiplied with the birth rate $\lambda$. Plugging $t=T-\tau$ in equation (5) gives the expected population size at time $\tau$ before sampling is proportional to $\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau}$.
The relative number of pendant branches of length $\tau$ per small time interval is thus given by
@@
The factor of two at the end of the first term accounts for the emergence of two branches of length $\tau$ when both offspring only have a single surviving descendant. $p_{>1}(\tau)=\sum_{n=2}^{\infty}p_{n}(\tau)$ denotes the probability that a clade founded at time $\tau$ has more than one extant member at the time of observation. Using equations (2) and (3), one obtains the probability density function of the lengths $\tau$ of pendant branches as
@@
For a finite time since the start of the birth-death process, $\tau$ has a finite support and hence $P_{\text{pend}}(\tau)$ cuts off to zero when $\tau$ is larger than the time $T$ since the start of the birth-death process. The normalizing factor of (10) is then found to be $(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T})^{2}/(\lambda(\lambda-(\lambda-2\mu)\operatorname{\mathrm{e}}^{-2(\lambda-\mu)T}-2\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T}))$. We focus on the particularly simple case of an asymptotically long running time $T\to\infty$. In this case, the cutoff moves to infinity and we find the normalizing factor in (10) to be one, making (10) a proper probability density function. We will focus on the asymptotic case $T\to\infty$ also when treating branches in the tree’s interior.
Curiously, distribution (10) coincides with the distribution of pendant branches conditioned on the population size $N$ (statistical scenario ii)), derived by Stadler and Steel Stadler and Steel (2012) as well as Mooers et al. Mooers et al. (2012). For details, see Appendix .1.
@@ IV Branch length distribution: interior branches above pendant branches
Figure 2: Statistics of pendant and interior branches. (A) This schematic phylogenetic tree shows birth events (circles) and death events ($\times$). The pendant branch of length $\tau$ on the left arose from a birth event in the past (solid circle), where one offspring (shown left) has exactly one extant descendant, and the other offspring (shown right) has at least one extant descendant. This particular pendant branch has one generation between its origin at time $\tau$ measured from the time of observation in the past and the present (not counting the birth event at $\tau$). (B) A birth event (top solid node) leading to a level-two branch of length $\tau_{2}$ and a level-one (pendant) branch of length $\tau_{1}$. (C) A schematic phylogenetic tree with branches of different levels, see text. (D) The distribution of branch lengths $\tau_{2}$ of second-level branches in the tree of a simulated population grown from a single individual over time $T=25$ as described in the text (grey bars), compared to the theoretical probability density function (11) in the asymptotic case of large $T$ marginalized over the pendant branch lengths $\tau_{1}$ (blue line). (E) The distribution of interior branch lengths in the same tree as before (grey bars), compared to the theoretical density function (14) (blue line).
Since the probability of a birth event at time $\tau_{1}+\tau_{2}$ was $\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)(\tau_{1}+\tau_{2})}$ (up to a normalizing prefactor, see below), the relative number of level-two and level-one branches of lengths $\tau_{2}$ and $\tau_{1}$ per small time intervals is thus
@@
Equation (11) gives the joint probability density function of branch lengths of pendant and level-two branches up to a normalization factor, which again turns out to be one. Figure 2D compares this result to numerical simulations. A population with birth rate $\lambda=1$ and death rate $\mu=0.75$ was grown from a single individual until time $T=25$ (where the population contained 3567 individuals), and the phylogenetic tree was reconstructed. The lengths of level-two branches were collected in a histogram and compared to the joint probability $P_{\text{level}}(\tau_{1},\tau_{2})$ given by (11) marginalized with respect to $\tau_{1}$.